OpenStudy (anonymous):

What is the y-intercept of the line represented by the equation below? y = –2x – 1 A. –2 B. –1 C. 1 D. 2

hartnn (hartnn):

y-intercept is when you plug in x= 0 in your equation

hartnn (hartnn):

y=- 2 (0) -1 =... ?

OpenStudy (anonymous):

y=0-1?

OpenStudy (anonymous):

-1?

hartnn (hartnn):

is correct :)

OpenStudy (anonymous):

thank you for the help :)

hartnn (hartnn):

welcome ^_^